OpenStudy (skylynnsmommy12):

type the correct answer in the box. use numerals instead of words. if necessary, use / for the fraction bar. the domain of the function f(x)= 3x^3 is {2,5} what is the function's range? the range of the function f(x)=3x^3 when its domain is {2,5} is_______?

OpenStudy (benlindquist):

@hayhayz

Vocaloid (vocaloid):

the domain {2,5} means that the only x-values included are x = 2 and x = 5 evaluate f(x) = 3x^3 when x = 2 and when x = 5 to get your range

OpenStudy (welshfella):



OpenStudy (welshfella):

= 3(2)^3 then repeat for x = 5

OpenStudy (pawanyadav):

Its an increasing function... So f(x) increase from 2 to 5 So its range is from f(2) to f(5).

OpenStudy (welshfella):

yes and f(2) means the value of the function when x = 2 in this case its 3(2)^3

OpenStudy (skylynnsmommy12):

thank yall.